OpenStudy (kiakat):

Claire starts driving with an initial velocity of 5.0 m/s towards north. She drives along a straight, level path with a uniform acceleration and travels 500. m in 20. seconds. A. What is the average velocity for the entire 20. seconds? B. What is the velocity at the end of 20. seconds?

OpenStudy (matt101):

When you want to calculate AVERAGE velocity in the presence of constant acceleration, you can safely use v=Δd/Δt to calculate your answer. In this case that will be 500/20=25 m/s. The question gives us initial velocity, distance, and time. To calculate the final velocity at 20 s, we can put together a couple kinematics equations: \[(1) \space \space v_f^2=v_i^2+2aΔ d\]\[(2) \space \space a = \frac{Δv}{Δt}\] Sub 2 back into 1 so that way we get rid of acceleration and only work with variables we know the value of to solve for final velocity: \[v_f^2=v_i^2+2 \frac{Δv}{Δt}Δd\] Note that Δv is final velocity minus initial velocity, Δt is just t, and Δd is just d. So: \[v_f^2=v_i^2+2 \frac{v_f-v_i}{t}d\]\[v_f^2t=v_i^2t+2dv_f-2dv_i\] I know this looks gross, but we can fix that - let's plug in the numbers we know: \[20v_f^2=500+1000v_f-5000\] Now put this in standard from and divide everything by 20 just to make things a bit easier: \[v_f^2-50v_f+225=0\] Solve this quadratic and you'll find that v=5 and v=45. We know that we're accelerating uniformly from 5 m/s, so our final speed can't also be 5 m/s. That means we can reject that value of v and go with 45 m/s as our final velocity! Whew! Let me know if you have any questions!